Ecology Question:

When Norway rats invade a new warehouse where living conditions are ideal, they multiply at a very rapid rate: r = 0.0157 per day. How many days are required for each doubling of the population size?

Select one:

a. Many days????

b. 2

c. 47.14

d. 44.14
Answer #1

td = 0.693 / g

td = doubling time and g = growth rate.

So, td = 0.693 / 0.0157 days = 44.14 days.

Hence the doubling time for population is 44.14 days.